Stuart's utility function for goods X and Y is represented as U(X,Y)=X0.8Y0.2. Assume that his income is $100 and the prices of goods X and Y are $20 and $10, respectively.

 


(d) Derive the demand curve for good X and demand curve for good Y.

 

Now a government subsidy program lowers the price of X from $20 per unit to $10 per unit.

 

(e) Calculate and graphically show the change in good X consumption resulting from the program.

 

(f) Graphically show the change in consumption attributable to the separate income and substitution effects.

 

(g) Show (graphically) how much the program cost the government.

 


1
Expert's answer
2021-12-05T18:58:32-0500

Solution:

d.). The demand curve for Good X:

Y = "0.25X\frac{Px}{Py}"

I = PxX + PyY

I = PxX + Py("0.25X\frac{Px}{Py}")

I = PxX + 0.25XPx

I = 1.25XPx

X = "\frac{I}{1.25Px}"

 

The demand curve for Good Y:

X = "4Y\frac{Py}{Px}"

I = PxX + PyY

I = Px("4Y\frac{Py}{Px}") + PyY

I = 4YPy + PyY

I = 5YPy

Y = "\frac{I}{5Py}"


e.). Budget constraint: I = PxX + PyY

100 = 20X + 10Y

New budget constraint after subsidy on Good X:

100 = 10X + 10Y

Good X = "\frac{100}{10}" = 10

Consumption of good X will increase by 5

This is depicted by the below graph:

 




 

f.). This is depicted by the below graph:



 

g.). This is depicted by the below graph: